# Distributional Limits for Eigenvalues of Graphon Kernel Matrices

### v3

@@ 1 Introduction
Spectral methods play a central role in the analysis of random graphs and network data Lei and Rinaldo (2015); Rohe et al. (2011). Their success in community detection Abbe (2018); Lei and Rinaldo (2015); Zhao et al. (2012), clustering Von Luxburg (2007), and goodness-of-fit procedures Bickel et al. (2011); Gao et al. (2015) has motivated a substantial body of work on the asymptotic behavior of eigenvalues and eigenvectors in latent-variable models. Classical results establish almost-sure or $L^{2}$ convergence of empirical kernel operators Koltchinskii and Giné (2000); El Karoui (2008); Braun (2006), together with consistency of eigenspaces and certain linear spectral statistics. However, the distributional behavior of individual eigenvalues, a question of central importance for uncertainty quantification remains comparatively underdeveloped.
@@
This paper develops a distributional theory for simple eigenvalues of kernel matrices arising from dense graphon-based random graphs. Working in the graphon framework of Lovász and Szegedy Lovász and Szegedy (2006), Lovász Lovász (2012), and Borgs et al. Borgs et al. (2008, 2017), we show that the fluctuations of each empirical eigenvalue exhibit a sharp probabilistic dichotomy. Depending on the degeneracy structure of the population eigenfunction, the leading fluctuations are either asymptotically Gaussian or converge to an explicit weighted chi-square law. A version of this dichotomy was first obtained by Chatterjee and Huang Chatterjee and Huang (2021) under Lipschitz regularity of the graphon. Our contribution is to extend this phenomenon to the minimal integrability setting $W\in L^{2}([0,1]^{2})\cap L^{\infty}([0,1]^{2})$, thereby covering discontinuous, non-smooth, and piecewise-constant models such as stochastic block models with unequal block sizes. The results demonstrate that smoothness assumptions, while technically convenient, are not intrinsic to the limiting behavior.
@@
From a probabilistic perspective, the fluctuation behavior is driven by the interaction between two sources of randomness: latent-position sampling and Bernoulli edge noise. A central finding of this work is that, at the $\sqrt{n}$ scale, the dominant contribution arises from the latent variables, while edge noise is asymptotically negligible. This separation, implicit in earlier work on spectral convergence Koltchinskii and Giné (2000); Zhu and Zhu (2024) and latent-position models Lei and Rinaldo (2015); Abbe (2018); Athreya et al. (2018), becomes explicit in our analysis and clarifies the sources of uncertainty in spectral procedures. In particular, the zero diagonal of the kernel matrix-an intrinsic structural feature of graphon models-plays a decisive role in the degenerate regime, where the linear Hoeffding projection vanishes and the limiting distribution becomes a weighted chi-square series.
@@
Our proof combines three probabilistic ingredients. First, a second-order Rayleigh-Schrödinger expansion for simple eigenvalues Katö (1995) reduces the problem to the analysis of a centered Rayleigh quotient. Second, concentration of kernel matrices under boundedness assumptions yields operator-norm control of the empirical perturbation. Third, a Hoeffding decomposition for the resulting symmetric statistic Serfling (2009) isolates the linear and degenerate components, revealing the mechanism behind the Gaussian versus weighted chi-square regimes. This synthesis demonstrates that the fluctuation dichotomy persists well beyond the smooth settings considered in Chatterjee and Huang (2021), and that the limiting distributions are determined by spectral structure and degeneracy rather than regularity of the kernel.
@@ 6 Proof of the main theorem
Moreover on this high-probability event, $\eta_{n}-\alpha_{n}\asymp n$, $\beta_{n}-\eta_{n}\asymp n$. The Kato-Temple inequality for self-adjoint matrices (see, e.g., Kato Katö (1995)) then yields
@@ Acknowledgments
This is a finite-dimensional specialization of the Kato-Temple inequality for self-adjoint operators; see  Katö (1995, Theorems V.4.10-V.4.11). ∎

### v4

@@ 1 Introduction
Spectral methods play a central role in the analysis of random graphs and network data Lei and Rinaldo (2015); Rohe et al. (2011). Their success in community detection Abbe (2017); Lei and Rinaldo (2015); Zhao et al. (2012), clustering Von Luxburg (2007), and goodness-of-fit procedures Bickel et al. (2011); Gao et al. (2015) has motivated a substantial body of work on the asymptotic behavior of eigenvalues and eigenvectors in latent-variable models. Classical results establish almost-sure or $L^{2}$ convergence of empirical kernel operators Koltchinskii and Giné (2000); Karoui (2008); Braun (2006), together with consistency of eigenspaces and certain linear spectral statistics. However, the distributional behavior of individual eigenvalues, a question of central importance for uncertainty quantification remains comparatively underdeveloped.
@@
This paper develops a distributional theory for simple eigenvalues of kernel matrices generated by dense graphon-based random graphs. Working within the graphon framework of Lovász and Szegedy Lovász and Szegedy (2006); Lovász (2012); Borgs et al. (2008, 2012), we show that each empirical eigenvalue exhibits a sharp probabilistic dichotomy in its fluctuations. Depending on the degeneracy structure of the population eigenfunction, the leading fluctuations are either asymptotically Gaussian or converge to an explicit weighted chi-square law. A version of this dichotomy was first obtained by Chatterjee and Huang (2025) under Lipschitz regularity of the graphon. Our contribution is to extend this phenomenon to the minimal integrability setting $W\in L^{2}([0,1]^{2})\cap L^{\infty}([0,1]^{2})$, thereby covering discontinuous, non-smooth, and piecewise-constant models such as stochastic block models with unequal block sizes. The results demonstrate that smoothness assumptions, while technically convenient, are not intrinsic to the limiting behavior.
@@
From a probabilistic perspective, the fluctuation behavior is driven by the interaction between two sources of randomness: latent-position sampling and Bernoulli edge noise. A central finding of this work is that, at the $\sqrt{n}$ scale, the dominant contribution arises from the latent variables, while edge noise is asymptotically negligible. This separation, implicit in earlier work on spectral convergence Koltchinskii and Giné (2000) and latent-position models Lei and Rinaldo (2015); Abbe (2017); Athreya et al. (2018), becomes explicit in our analysis and clarifies the sources of uncertainty in spectral procedures. In particular, the zero diagonal of the kernel matrix-an intrinsic structural feature of graphon models-plays a decisive role in the degenerate regime, where the linear Hoeffding projection vanishes and the limiting distribution becomes a weighted chi-square series.
@@
Our proof combines three probabilistic ingredients. First, a second-order Rayleigh-Schrödinger expansion for simple eigenvalues Kato and Katåo (1966) reduces the problem to the analysis of a centered Rayleigh quotient. Second, concentration of kernel matrices under boundedness assumptions yields operator-norm control of the empirical perturbation. Third, a Hoeffding decomposition for the resulting symmetric statistic Serfling (2009) isolates the linear and degenerate components, revealing the mechanism behind the Gaussian versus weighted chi-square regimes. This synthesis demonstrates that the fluctuation dichotomy persists well beyond the smooth settings considered in Chatterjee and Huang (2025), and that the limiting distributions are determined by spectral structure and degeneracy rather than regularity of the kernel.
@@ 6 Proof of the main theorem
Moreover on this high-probability event, $\eta_{n}-\alpha_{n}\asymp n$, $\beta_{n}-\eta_{n}\asymp n$. The Kato-Temple inequality for self-adjoint matrices (see, e.g., Kato Kato and Katåo (1966)) then yields
@@ Acknowledgments
This is a finite-dimensional specialization of the Kato-Temple inequality for self-adjoint operators; see  Kato and Katåo (1966, Theorems V.4.10-V.4.11). ∎